f(x)=4x+2 and g(x)=3x-1, find (f-g)(4)






1
Expert's answer
2021-11-09T16:48:41-0500

f(x) = 4x + 2

g(x) = 3x - 1

(f - g)(4) = 4(4) + 2 - 3(4) - 1

= 16 + 2 - 12 - 1

= 18 - 12 - 1

= 6 - 1

(f - g)(4) = 5